# Review: Quantum Metrology and Sensing with Many-Body Systems

### v2

@@ X Non-Equilibrium Quantum Sensing: Floquet Phase Transition
Quantum metrologically advantageous states can also be generated using the application of external time-periodic driving to the system. In this case, the frequency of the external driving can resonate with the lowest energies of the unperturbed systems. This resonance can be harnessed for quantum enhanced sensing Lang et al. (2015); Wang et al. (2022b); Jiang et al. (2022); Pang and Jordan (2017); Zhou et al. (2020); Mishra and Bayat (2021, 2022); Yu et al. (2022); Ivanov (2021). This resonance phenomena has also been observed in gaped many-body quantum systems when the frequency of external driving matches with the lowest energy gap of the unperturbed Hamiltonian. Russomanno et al. (2016); Mishra and Bayat (2021, 2022). Various important physical features of systems with time-periodic influence can be extracted by employing Floquet formalism. For a time-periodic Hamiltonian, $H(t+\tau)=H(t)$, with periodicity $\tau=2\pi/\omega$, the solution of the Schrödinger equation follows from the Floquet theorem. The Floquet theorem gives an ansatz of the form $|\Psi(t)\rangle=\sum_{\ell}e^{-i\mu_{\ell}t}|\Phi_{\ell}(t)\rangle$. Here $\mu_{\ell}^{\prime}s$ are eigenphase (the Floquet phase), $|\Phi_{\ell}(t)\rangle$ the Floquet modes, and $\ell\in\mathbb{Z}$. Substituting the ansatz to the Schrödinger equation, $i\frac{\partial|\Psi(t)\rangle}{\partial t}=H(t)|\Psi(t)\rangle$, the Floquet modes satisfy
$\Big{(}H(t)-i\frac{\partial}{\partial t}\Big{)}|\Phi_{\ell}(t)\rangle=\mu_{\ell}|\Phi_{\ell}(t)\rangle.$ (70)
It can be noted that $|\Phi_{\ell}(t+\tau)\rangle$ is also a solution of the above equation with Floquet eigenphase $\mu_{\ell}$, therefore, one can write $|\Phi_{\ell}(t+\tau)\rangle=|\Phi_{\ell}(t)\rangle$. Furthermore, the time evolution propagator, describing the dynamics of the system, is given by
@@
where ${\cal T}_{t}$ represent time ordering operator. It is noted that $U(t+\tau)|\Psi(t)\rangle=|\Psi(t+\tau)\rangle$. Using the ansatz for $|\Psi(t)\rangle$ in the above equation, one gets $U(t+\tau)e^{-i\mu_{\ell}t}|\Phi_{\ell}(t)\rangle=e^{-i\mu_{\ell}t}|\Phi_{\ell}(t+\tau)\rangle$. This shows that the Floquet modes are eigenvectors of the one time-period propagator $U(\tau,0)$ with $\{e^{-i\mu_{\ell}\tau}\}$ the corresponding eigenvalues.
@@
Properties of Floquet modes $\{|\Phi_{\ell}(t)\}$ and Floquet phases $\{\mu_{\ell}\}$ have been explored for sensing periodic magnetic fields due to clusters of few qubits Lang et al. (2015). Here, the effective magnetic field due to the cluster of spins is detected using a single qubit as a sensor. The initial state of the system is $|\psi(0)\rangle{=}\frac{1}{\sqrt{2}}(\ket{\uparrow}{+}\ket{\downarrow})\otimes|\mathcal{B}(0)\rangle$, where |$\ket{\mathcal{B}(0)}$ is the initial state of the detected spin cluster at time $t=0$ and $\ket{\uparrow}$ , $\ket{\downarrow}$ are the computational basis states of the electronic spin, respectively. At later times, due to the interaction with the sensor (the qubit state), the spin cluster state becomes entangled with the sensor. The magnetic field signal is detected from the data of temporal coherence of the time-evolved state of the detected spin cluster. The coherence can be expressed in terms of the eigenvalues and eigenstates of the one-period unitary evolution operator or Floquet operator. The coherence displays a dip which occurs at the avoided crossings of the Floquet eigenstates where $e^{i\mu_{\ell}}\approx e^{-i\mu_{k}}$ or when the Floquet gap, $\Delta_{F}=\mu_{\ell}-\mu_{k}$, vanishes. Thus, by measuring the location of the coherence dip, the energy-difference between Floquet states can be observed with precision higher than the one predicted by the average Hamiltonian theory. However, this method can only be applicable for a sensing around the resonance frequencies. Sensing of arbitrary frequency fields has been proposed using a single qubit as a sensor in Floquet systems Wang et al. (2022b). In addition, periodically driven systems and the transition of Floquet eigenphases have been used for magnetic-field signal amplification Jiang et al. (2022). In Ref. Pang and Jordan (2017), an optimal control method is developed to estimate the amplitude, $B$, of a time-periodic magnetic field $B(t)=B(\cos(\omega t)\sigma_{x}+\sin(\omega t)\sigma_{z})$. In Ref. Yang et al. (2022b), a systematic variational approach for establishing optimal control was set up, where it was shown that for magnetometry, restricted controls in the form of one and two-body interactions are already sufficient to approximately achieve Heisenberg scaling. It is shown here that the QFI with respect to $B$ can achieve super-Heisenberg scaling with interrogation time, $t$, by applying an optimal control pulse as compared to the one without control. A solid state sensor with an ensemble of high density Nitrogen Vacancy (NV) centers in diamonds has been investigated both theoretically and experimentally by Floquet formalism  Zhou et al. (2020). The system, here, is a dense ensemble of electronic spins. Each electronic spin act as an independent probe for the external signal. Thus, one would expect that, by increasing the number of particles in the system, the sensitivity can be improved. The improvement, however, is severely hindered by interaction between the spins, since a large spins are confined in a small volume, the spin-spin interaction is unavoidable. By applying a series of dynamical decoupling pulses, the interaction between the spins has been suppressed, which results in a larger coherence time. In this way the standard quantum limit is shown to be restored which is lost due to spin-spin interaction.
@@
It is important to note that the above sensing protocols either do not employ interaction between the particles directly or use schemes to suppress the particle-particle interaction. Thus, the above schemes are equivalent to the single-particle probe quantum metrology schemes with several rounds of repetition of the protocol. It is natural to extend the formalism to many-body systems with Floquet dynamics without applying any dynamical decoupling pulses and ask if still one can harness any property for improving the precision well beyond the standard limit. Sensing protocols have been developed in many-body systems using Floquet formalism for detecting both static Mishra and Bayat (2021) and periodic fields Mishra and Bayat (2022). The Hamiltonian considered in those studies is
$H(t)$ $=$ $-\frac{J}{2}\sum_{i=1}^{N}\Big{[}\Big{(}\frac{1+\gamma}{2}\Big{)}\sigma^{x}_{i}\sigma^{x}_{i+1}+\Big{(}\frac{1-\gamma}{2}\Big{)}\sigma^{y}_{i}\sigma^{y}_{i+1}\Big{]}$ (72)
-- $\frac{(h_{0}+h(t))}{2}\sum_{i}\sigma^{z}_{i},$
@@
where, $\sigma^{\alpha}$ ($\alpha=x,y,z$) are the Pauli matrices, $J$ is the exchange coupling, $-1{\leq}\gamma{\leq}1$ is the anisotropic parameter, and the periodic-boundary conditions, i.e., $\sigma^{\alpha}_{N+1}{\equiv}\sigma^{\alpha}_{1}$, is imposed. Furthermore, the time-dependent field $h(t)$ is periodic with a period $\tau=\frac{2\pi}{\omega}$. Thus, the Hamiltonian is a time-periodic $H(t+n\tau)=H(t)$, where $n\in\mathbb{Z}$. In the absence of the periodic field, i.e. $h(t)=0$, the Hamiltonian (72) shows quantum criticality at $h_{0}{=}h_{c}{=}J$ for all values of $\gamma$ Sachdev and Safari (2011). Using Floquet formalism for solving the dynamics driven by the Hamiltonian in Eq. (72), one can estimate both the static field $h_{0}$ Mishra and Bayat (2021) and the amplitude of the periodic field $h(t)$ Mishra and Bayat (2022) even with partial accessibility to a block of spins. Note that partial accessibility implies that one only has access to a certain block of spins to perform measurement and gains information about the parameter of interest. It is to be emphasis here that the sensitivity to the parameter of interest naturally deceases as the accessibility reduces to smaller blocks. For example, it is shown that the QFI with respect to $h_{0}$ in the ground state of Eq. (72), when $h(t)=0$, decreases as compared to the full system Mishra and Bayat (2021). By driving the system periodically using an external drive field $h(t)$ the QFI can be made to exhibits peaks not only at the critical field, but at the points where the Floquet resonances occur. Thus, even with partial accessibility, the probe state becomes more sensitive with respect to the magnetic field $h_{0}$ when a suitable controlled driving magnetic field is applied such that the floquet gap closes.
@@
$U(\tau,0)=e^{-iH_{F}\tau},$ (73)
where $H_{F}$ is called the Floquet Hamiltonian and it describes the evolution of the system at the stroboscopic times $t=n\tau$. The knowledge of the state of the system at the later times $t=n\tau$ can be fully obtained by successive application of $U(\tau,0)$. The $U(\tau,0)$ admits spectral decomposition in terms of eigenvectors $\{|\mu\rangle\}$ and eigenvalues $\{\mu(\tau)\}$ of $U(\tau,0)$ which depends on unknown parameter to be estimated. The eigenvalues are periodic and are given by $\mu_{\ell}=\mu_{\ell}+n\omega$. Starting from a initial state, $|\Psi_{0}\rangle$, the time-evolved state can be written as $|\Psi(n\tau)\rangle{=}\sum_{i}e^{-{\it i}\mu_{i}n\tau}|\mu_{i}\rangle|\langle\mu_{i}|\Psi_{0}\rangle|^{2}$.
@@
Now for periodic field sensing, the Hamiltonian in Eq. (72) is taken to be a a transverse -field Ising model ($\gamma=1$) with a time-dependent magnetic field of unknown amplitude $h_{1}$ added to the Hamiltonian Mishra and Bayat (2022). For simplicity, the form is taken as Dirac-delta pulse $h(t)=h_{1}\sum_{n=0}^{\infty}\delta(t-n\tau)$. The aim is to estimate the unknown amplitude $h_{1}$. The $I^{ss}_{Q}$ is maximum and has a peak along a line in the $(h_{0}-h_{1})$ plane. Interestingly, this line corresponds to the line of vanishing Floquet gap. Moreover, in the steady-state, the Fisher information $I^{ss}_{Q}$ with respect to $h_{1}$ exhibits the Heisenberg type scaling $I^{ss}_{Q}\sim L^{\beta}$. An important finding is that at the location of the vanishing Floquet gap, the exponent $\beta\approx 1.96$ while for points where the Floquet phases do not cross, the exponent decreases towards the standard limit.
@@ XI Non-Equilibrium Quantum Sensing: Dissipative Phase Transitions
Physical systems unavoidably interact with their environment, often referred to as the bath or reservoir, which possesses a larger number of degrees of freedom than the system of interest. The study of open quantum systems, which involves understanding the interaction between our system and the environment, has led to remarkable theoretical findings and practical applications Breuer et al. (2002). However, the presence of noisy environments causes the systems to lose their coherence (decohere) and, consequently, results in unavoidable detrimental effects on their dynamics. This issue is particularly relevant for upcoming quantum technologies, as fighting against decoherence is essential for achieving more efficient quantum probes Goldstein et al. (2011), developing algorithms for specific quantum tasks Plenio et al. (1997), and quantum simulators with significantly longer coherence times Kjaergaard et al. (2020). Given that the interaction between the relevant system and its environment is inevitable, one might wonder whether it is possible to overcome or harness the detrimental effects of decoherence.
@@ XI.1 Profiting from dissipation
Since non-classical states are crucial for countless applications and theoretical advances, it would be highly desirable to have a mechanism where instead of avoiding dissipation, one could harness the very process of dissipation to one’s advantage. The seminal work by Alfred Kastler in 1950 pointed in that direction Kastler, Alfred (1950). The process, known as optical pumping, prepares any initial quantum state into a well-defined target state via selection rules of spontaneous emission. Building on Kastler’s work, coherent population trapping was experimentally demonstrated in the 1970s by Arimondo and Orriols Alzetta et al. (1976); Arimondo and Orriols (1976). In a nutshell, a three-level energy configuration with transitions levels $|a\rangle\longleftrightarrow|b\rangle$ and $|a\rangle\longleftrightarrow|c\rangle$ and Hamiltonian $H=\hbar\Omega(|a\rangle\langle b|-|a\rangle\langle c|+h.c)$ allows for a quantum state with a zero eigenvalue (a dark state). Remarkably, such a dark state, denoted as $|\psi_{d}\rangle\sim|b\rangle+|c\rangle$, decouples from the external pumping field $\Omega$, thus allowing excitations that reach that energy level to remain there. In contrast, the bright state, denoted as $|\psi_{b}\rangle\sim|b\rangle-|c\rangle$, oscillates with the driving field $\Omega$ and the state $|a\rangle$. Indeed, this idea lies at the heart of laser cooling Aspect et al. (1988) as the ground state represents another specific quantum state. The paradigm shift of harnessing dissipation as a resource rather than a drawback has led to systematic studies of decoherence itself Myatt et al. (2000), the ability to prepare a quantum state from any arbitrary initial state Poyatos et al. (1996), the use of atomic engineered reservoirs to create two-mode Einstein-Podolsky-Rosen-entangled states Pielawa et al. (2007, 2010), protection against noise through the engineering of pointer states Carvalho et al. (2001), dissipative quantum computation Verstraete et al. (2009), the creation of many-body states and non-equilibrium quantum phases Diehl et al. (2008), thermometry of cold Fermi gases with dephasing-impurities Mitchison et al. (2020); Bouton et al. (2020); Brattegard and Mitchison (2024), quantum memories using dissipative mechanisms Pastawski et al. (2011), the generation of Bell states on an open-system quantum simulator Barreiro et al. (2011), and the possibility to enhance sensing precision nearly to the Heisenberg limit Goldstein et al. (2011).

### v3

@@ X Non-Equilibrium Quantum Sensing: Floquet Phase Transition
Quantum metrologically advantageous states can also be generated using the application of external time-periodic driving to the system. In fact, in interferometry-based quantum sensing one can overcome the detrimental effect of decoherence by driving the system in a proper way to restore Heisenberg scaling Bai and An (2023). In periodically driven systems, the frequency of the external driving can resonate with the lowest energies of the unperturbed systems. This resonance can be harnessed for quantum-enhanced sensing Lang et al. (2015); Wang et al. (2022b); Jiang et al. (2022); Pang and Jordan (2017); Zhou et al. (2020); Mishra and Bayat (2021, 2022); Yu et al. (2022); Ivanov (2021). This resonance phenomena has also been observed in gaped many-body quantum systems when the frequency of external driving matches with the lowest energy gap of the unperturbed Hamiltonian. Russomanno et al. (2016); Mishra and Bayat (2021, 2022). Various important physical features of systems with time-periodic influence can be extracted by employing Floquet formalism. For a time-periodic Hamiltonian, $\hat{H}(t+\tau)=\hat{H}(t)$, with periodicity $\tau=2\pi/\omega$, the solution of the Schrödinger equation follows from the Floquet theorem. The Floquet theorem gives an ansatz of the form $|\Psi(t)\rangle=\sum_{\ell}e^{-i\mu_{\ell}t}|\Phi_{\ell}(t)\rangle$. Here $\mu_{\ell}^{\prime}s$ are quasienergies (the Floquet phase), $|\Phi_{\ell}(t)\rangle$ the Floquet modes, and $\ell\in\mathbb{Z}$. Substituting the ansatz to the Schrödinger equation, $i\frac{\partial|\Psi(t)\rangle}{\partial t}=\hat{H}(t)|\Psi(t)\rangle$, the Floquet modes satisfy
$\Big{(}\hat{H}(t)-i\frac{\partial}{\partial t}\Big{)}|\Phi_{\ell}(t)\rangle=\mu_{\ell}|\Phi_{\ell}(t)\rangle.$ (72)
It can be noted that $|\Phi_{\ell}(t+\tau)\rangle$ is also a solution of the above equation with Floquet quasienergies $\mu_{\ell}$, therefore, one can write $|\Phi_{\ell}(t+\tau)\rangle=|\Phi_{\ell}(t)\rangle$. Furthermore, the time evolution propagator, describing the dynamics of the system, is given by
@@
where ${\cal T}_{t}$ represent time ordering operator. It is noted that $\hat{U}(t+\tau)|\Psi(t)\rangle=|\Psi(t+\tau)\rangle$. Using the ansatz for $|\Psi(t)\rangle$ in the above equation, one gets $\hat{U}(t+\tau)e^{-i\mu_{\ell}t}|\Phi_{\ell}(t)\rangle=e^{-i\mu_{\ell}t}|\Phi_{\ell}(t+\tau)\rangle$. This shows that the Floquet modes are eigenvectors of the one time-period propagator $\hat{U}(\tau,0)$ with $\{e^{-i\mu_{\ell}\tau}\}$ the corresponding eigenvalues.
@@
Properties of Floquet modes $\{|\Phi_{\ell}(t)\}$ and Floquet phases $\{\mu_{\ell}\}$ have been explored for sensing periodic magnetic fields due to clusters of few qubits Lang et al. (2015). Here, the effective magnetic field due to the cluster of spins is detected using a single qubit as a sensor. The initial state of the system is $|\psi(0)\rangle{=}\frac{1}{\sqrt{2}}(\ket{\uparrow}{+}\ket{\downarrow})\otimes|\mathcal{B}(0)\rangle$, where |$\ket{\mathcal{B}(0)}$ is the initial state of the detected spin cluster at time $t=0$ and $\ket{\uparrow}$ , $\ket{\downarrow}$ are the computational basis states of the electronic spin, respectively. At later times, due to the interaction with the sensor (the qubit state), the spin cluster state becomes entangled with the sensor. The magnetic field signal is detected from the data of temporal coherence of the time-evolved state of the detected spin cluster. The coherence can be expressed in terms of the eigenvalues and eigenstates of the one-period unitary evolution operator or Floquet operator. The coherence displays a dip which occurs at the avoided crossings of the Floquet eigenstates where $e^{i\mu_{\ell}}\approx e^{-i\mu_{k}}$ or when the Floquet gap, $\Delta_{F}=\mu_{\ell}-\mu_{k}$, vanishes. Thus, by measuring the location of the coherence dip, the energy-difference between Floquet states can be observed with precision higher than the one predicted by the average Hamiltonian theory. However, this method can only be applicable for a sensing around the resonance frequencies. Sensing of arbitrary frequency fields has been proposed using a single qubit as a sensor in Floquet systems Wang et al. (2022b). In addition, periodically driven systems and the transition of Floquet quasienergies have been used for magnetic-field signal amplification Jiang et al. (2022). In Ref. Pang and Jordan (2017), an optimal control method is developed to estimate the amplitude, $B$, of a time-periodic magnetic field $B(t)=B(\cos(\omega t)\hat{\sigma}_{x}+\sin(\omega t)\hat{\sigma}_{z})$. In Ref. Yang et al. (2022b), a systematic variational approach for establishing optimal control was set up, where it was shown that for magnetometry, restricted controls in the form of one and two-body interactions are already sufficient to approximately achieve Heisenberg scaling. It is shown here that the QFI with respect to $B$ can achieve quantum-enhanced sensitivity with interrogation time, $t$, by applying an optimal control pulse as compared to the one without control. A solid state sensor with an ensemble of high density Nitrogen Vacancy (NV) centers in diamonds has been investigated both theoretically and experimentally by Floquet formalism  Zhou et al. (2020). The system, here, is a dense ensemble of electronic spins. Each electronic spin act as an independent probe for the external signal. Thus, one would expect that, by increasing the number of particles in the system, the sensitivity can be improved. The improvement, however, is severely hindered by interaction between the spins, since a large spins are confined in a small volume, the spin-spin interaction is unavoidable. By applying a series of dynamical decoupling pulses, the interaction between the spins has been suppressed, which results in a larger coherence time. In this way the standard quantum limit is shown to be restored which is lost due to spin-spin interaction.
@@
It is important to note that the above sensing protocols either do not employ interaction between the particles directly or use schemes to suppress the particle-particle interaction. Thus, the above schemes are equivalent to the single-particle probe quantum metrology schemes with several rounds of repetition of the protocol. It is natural to extend the formalism to many-body systems with Floquet dynamics without applying any dynamical decoupling pulses and ask if still one can harness any property for improving the precision well beyond the standard quantum limit. Sensing protocols have been developed in many-body systems using Floquet formalism for detecting both static Mishra and Bayat (2021) and periodic fields Mishra and Bayat (2022). The Hamiltonian considered in those studies is
$\hat{H}(t)$ $=$ $-\frac{J}{2}\sum_{i=1}^{N}\Big{[}\Big{(}\frac{1+\gamma}{2}\Big{)}\hat{\sigma}^{x}_{i}\hat{\sigma}^{x}_{i+1}+\Big{(}\frac{1-\gamma}{2}\Big{)}\hat{\sigma}^{y}_{i}\hat{\sigma}^{y}_{i+1}\Big{]}$ (74)
-- $\frac{(h_{0}+h(t))}{2}\sum_{i}\hat{\sigma}^{z}_{i},$
@@
where, $\hat{\sigma}^{x,y,z}$ are the Pauli matrices, $J$ is the exchange coupling, $-1{\leq}\gamma{\leq}1$ is the anisotropic parameter, and the periodic-boundary conditions, i.e., $\hat{\sigma}^{x,y,z}_{N+1}{\equiv}\hat{\sigma}^{x,y,z}_{1}$, is imposed. Furthermore, the time-dependent field $h(t)$ is periodic with a period $\tau=\frac{2\pi}{\omega}$. Thus, the Hamiltonian is a time-periodic $\hat{H}(t+n\tau)=\hat{H}(t)$, where $n\in\mathbb{Z}$. In the absence of the periodic field, i.e. $h(t)=0$, the Hamiltonian (74) shows quantum criticality at $h_{0}{=}h_{c}{=}J$ for all values of $\gamma$ Sachdev (1999). Using Floquet formalism for solving the dynamics driven by the Hamiltonian in Eq. (74), one can estimate both the static field $h_{0}$ Mishra and Bayat (2021) and the amplitude of the periodic field $h(t)$ Mishra and Bayat (2022) even with partial accessibility to a block of spins. Note that partial accessibility implies that one only has access to a certain block of spins to perform measurement and gains information about the parameter of interest. It is to be emphasis here that the sensitivity to the parameter of interest naturally deceases as the accessibility reduces to smaller blocks. For example, it is shown that the QFI with respect to $h_{0}$ in the ground state of Eq. (74), when $h(t)=0$, decreases as compared to the full system Mishra and Bayat (2021). By driving the system periodically using an external drive field $h(t)$ the QFI can be made to exhibits peaks not only at the critical field, but at the points where the Floquet resonances occur. Thus, even with partial accessibility, the probe state becomes more sensitive with respect to the magnetic field $h_{0}$ when a suitable controlled driving magnetic field is applied such that the floquet gap closes.
@@
$\hat{U}(\tau,0)=e^{-i\hat{H}_{F}\tau},$ (75)
where $\hat{H}_{F}$ is called the Floquet Hamiltonian and it describes the evolution of the system at the stroboscopic times $t=n\tau$. The knowledge of the state of the system at the later times $t=n\tau$ can be fully obtained by successive application of $\hat{U}(\tau,0)$. The $\hat{U}(\tau,0)$ admits spectral decomposition in terms of eigenvectors $\{|\mu\rangle\}$ and eigenvalues $\{\mu(\tau)\}$ of $\hat{U}(\tau,0)$ which depends on unknown parameter to be estimated. The eigenvalues are periodic and are given by $\mu_{\ell}=\mu_{\ell}+n\omega$. Starting from a initial state, $|\Psi_{0}\rangle$, the time-evolved state can be written as $|\Psi(n\tau)\rangle{=}\sum_{i}e^{-{\it i}\mu_{i}n\tau}|\mu_{i}\rangle|\langle\mu_{i}|\Psi_{0}\rangle|^{2}$.
@@
Now for periodic field sensing, the Hamiltonian in Eq. (74) is taken to be a a transverse -field Ising model ($\gamma=1$) with a time-dependent magnetic field of unknown amplitude $h_{1}$ added to the Hamiltonian Mishra and Bayat (2022). For simplicity, the form is taken as Dirac-delta pulse $h(t)=h_{1}\sum_{n=0}^{\infty}\delta(t-n\tau)$. The aim is to estimate the unknown amplitude $h_{1}$. The $I^{ss}_{Q}$ is maximum and has a peak along a line in the $(h_{0}-h_{1})$ plane. Interestingly, this line corresponds to the line of vanishing Floquet gap. Moreover, in the steady-state, the Fisher information $I^{ss}_{Q}$ with respect to $h_{1}$ exhibits the Heisenberg type scaling $I^{ss}_{Q}\sim L^{\beta}$. An important finding is that at the location of the vanishing Floquet gap, the exponent $\beta\approx 1.96$ while for points where the Floquet phases do not cross, the exponent decreases towards the standard quantum limit.
@@ XI Non-Equilibrium Quantum Sensing: Dissipative Phase Transitions
Physical systems unavoidably interact with their environment, often referred to as the bath or reservoir, which possesses a larger number of degrees of freedom than the system of interest. The study of open quantum systems, which involves understanding the interaction between our system and the environment, has led to remarkable theoretical findings and practical applications Breuer et al. (2002). However, the presence of noisy environments causes the systems to lose their coherence (decohere) and, consequently, results in unavoidable detrimental effects on their dynamics. This issue is particularly relevant for upcoming quantum technologies, as fighting against decoherence is essential for achieving more efficient quantum probes Goldstein et al. (2011), developing algorithms for specific quantum tasks Plenio et al. (1997), and quantum simulators with significantly longer coherence times Kjaergaard et al. (2020). See Ref. Jiao et al. (2023) for a recent and comprehensive review on quantum metrology in the noisy intermediate-scale quantum (NISQ) era. Moreover, various sensing strategies have been proposed to estimate unknown parameters in open quantum systems. A few notable examples include recent works Wang et al. (2017a); Bai et al. (2019); Wu et al. (2021b, c); Wu and An (2021)—see also Ref. Mehboudi et al. (2019) for a thorough and comprehensive review on thermometry. Since the interaction between the system and its environment is unavoidable, a natural question arises: Can the detrimental effects of decoherence be mitigated or even exploited to enhance metrological performance? In the following, we review these possibilities in detail.
@@ XI.1 Profiting from dissipation
Since non-classical states are crucial for countless applications and theoretical advances, it would be highly desirable to have a mechanism where instead of avoiding dissipation, one could harness the very process of dissipation to one’s advantage. The seminal work by Alfred Kastler in 1950 pointed in that direction Kastler, Alfred (1950). The process, known as optical pumping, prepares any initial quantum state into a well-defined target state via selection rules of spontaneous emission. Building on Kastler’s work, coherent population trapping was experimentally demonstrated in the 1970s by Arimondo and Orriols Alzetta et al. (1976); Arimondo and Orriols (1976). In a nutshell, a three-level energy configuration with transitions levels $|a\rangle\longleftrightarrow|b\rangle$ and $|a\rangle\longleftrightarrow|c\rangle$ and Hamiltonian $\hat{H}=\hbar\Omega(|a\rangle\langle b|-|a\rangle\langle c|+h.c)$ allows for a quantum state with a zero eigenvalue (a dark state). Remarkably, such a dark state, denoted as $|\psi_{d}\rangle\sim|b\rangle+|c\rangle$, decouples from the external pumping field $\Omega$, thus allowing excitations that reach that energy level to remain there. In contrast, the bright state, denoted as $|\psi_{b}\rangle\sim|b\rangle-|c\rangle$, oscillates with the driving field $\Omega$ and the state $|a\rangle$. Indeed, this idea lies at the heart of laser cooling Aspect et al. (1988) as the ground state represents another specific quantum state. The paradigm shift of harnessing dissipation as a resource rather than a drawback has led to systematic studies of decoherence itself Myatt et al. (2000), the ability to prepare a quantum state from any arbitrary initial state Poyatos et al. (1996), the use of atomic engineered reservoirs to create two-mode Einstein-Podolsky-Rosen-entangled states Pielawa et al. (2007, 2010), protection against noise through the engineering of pointer states Carvalho et al. (2001), dissipative quantum computation Verstraete et al. (2009), the creation of many-body states and non-equilibrium quantum phases Diehl et al. (2008), thermometry of cold Fermi gases with dephasing-impurities Mitchison et al. (2020); Bouton et al. (2020); Brattegard and Mitchison (2024), quantum memories using dissipative mechanisms Pastawski et al. (2011), the generation of Bell states on an open-system quantum simulator Barreiro et al. (2011), and the possibility to enhance sensing precision nearly to the Heisenberg limit Goldstein et al. (2011).